Grade upto college level Physical Chemistry

A metal crystallizes into cubic phases, face centered cubic (BCC), whose unit cell lengths are 3.5 and 3.0 Å, respectively, Calculate the ratio of densities of FCC and BCC.

Shane Macguire
12 Years agoGrade upto college level

1 Answer

Deepak Patra
12 Years ago
Hello Student,
Please find the answer to your question
Density in fcc = n1 * at.wt./V1 *No.
Density in bcc = n2 * at.wt./V2 *No.
fcc unit cell length = 3.5 Å
bcc unit cell length = 3.0 Å
Density in fcc = n1 * at.wt./V1 *Av.No.
Density in bcc = n2 * at.wt./V2 *Av.No.
∴ Dfcc/Dbcc = n1/n2 * V2/V1
n1 for fcc = 4; Also V1 = a3 = (3.5 * 10-8)3
n2 for fcc = 2; Also V2 = a3 = (3.0 * 10-8)3
∴ Dfcc/Dbcc = 4 * (3.0 * 10-8)3/2 * (3.5 * 10-8)3 = 1.259

Thanks
Deepak patra
askIITians Faculty